Before Ronald sold two female rabbits, half of his rabbits were female. After the sale, only one third of his rabbits were female. If x represents his original number of rabbits then,
1/2 x -2=1/3(x-2)
Solve this equation to find the number of rabbits that he had before the sale.

Answer by stanbon(75887)

1/2 x -2=1/3(x-2)
Solve this equation
--------------
Multiply both sides by 6 to get:
3x - 12 = 2(x-2)
3x-12 = 2x-4
x = 8
==============
Cheers,
Stan H.